what is the answer for the problem slope=4, passing through (1,3)
Answer by tutorcecilia(2144)

slope=4, passing through (1,3)[Use the point-intercept formula]
(y-y1)=m(x-x1) [Plug-in the values]
(3-y)=4(1-x) [Simplify]
3-y=4-4x [Solve for y]
3-3-y=4-4x-3
-y=-4x+4-3
-y=-4x+1
-y/-1=-4x/-1 + 1/-1
y=4x-1